OpenStudy (anonymous):

find d^2y/dx^2 in terms of x and y. x^2-y^2=16

OpenStudy (anonymous):

this is a calculus problem

OpenStudy (allank):

You can use implicit differentiation..can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats what i"m trying to do but I don't get the right answer -16/y^3

OpenStudy (allank):

Nice so we will first do d/dx for the whole function to get: 2x-2y*(dy/dx)=0 Are you okay with this first one?

OpenStudy (anonymous):

yes

OpenStudy (allank):

Then we will do the second d/dx for the whole function to get: 2-(2\frac{dy}{dx}+2y*\frac{d2y}{dx2})=0

OpenStudy (allank):

Sorry, that didn't come out right...

OpenStudy (allank):

You'll have: \[2-(2\frac{dy}{dx}\frac{dy}{dx}+2y\frac{d2y}{dx2})\]

OpenStudy (allank):

And that =0 of course.

OpenStudy (anonymous):

why do you take the 2 out?

OpenStudy (allank):

I had this: 2x-2y*(dy/dx)=0 Then d/dx of 2x is 2, and the second part is treated as one block covered by that negative sign.

OpenStudy (allank):

Or are you asking about the 2 within the parentheses?

OpenStudy (anonymous):

no I got it 2 is the derivative of 2x

OpenStudy (allank):

Good. Can you take it from there?

OpenStudy (anonymous):

no I can't

OpenStudy (anonymous):

are you suppose to cancel out the derivatives?

OpenStudy (allank):

No, what you want to do is to use algebra to make d2y/dx2 the subject of the equation. Leave the derivatives intact.

OpenStudy (allank):

That should give you \[\frac{d2y}{dx2}=\frac{2-2\frac{dy}{dx}\frac{dy}{dx}}{2y} \]

OpenStudy (allank):

I have to be honest, I'm not entirely sure whether we should leave the \[\frac{dy}{dx}\frac{dy}{dx} \ \] in that form.

OpenStudy (anonymous):

I think its just dy/dx

OpenStudy (allank):

I'm more inclined to leave it as (dy/dx)^2 but you can confirm from other openstudy users. What do you think, @radar ?

OpenStudy (radar):

Sorry but I can not offer any clarification. Hopefully another OC user can, or repost the question.